penicillin -n is an antibacterial agent that contains 8.8% sulphur  by mass. what could be the minimum molar of penicilin-n??


plzzzzz........... explain.............

Grade:Upto college level

1 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Dear student,

let there are m number of sulphur atoms

=> 8.8*M/100 = m*32           (M is molecular mass of pencilin)

=> M = 3200/8.8 *m = 4000/11 *m

for m= 1 M = 363.63

Minimum mola mass of pencilin = 363.63